# Black Hole Quasi-Periodic Oscillations in the Presence of Gauss-Bonnet Trace Anomaly

### v1

@@ Abstract
We investigate the effects of the Gauss-Bonnet (GB) gravitational trace anomaly on the circular motion of test particles around black holes (BHs) and its implications for quasi-periodic oscillations (QPOs) in various theoretical models. Beginning with the equations of motion, we study the effects on effective potential, angular momentum, specific energy, and the innermost stable circular orbit (ISCO) induced by the anomaly parameter $\alpha$. The fundamental frequencies are calculated. Moreover, we examine several QPO models, including PR, RP, WD, TD, and ER2–ER4, and study the relationship between the upper and lower QPO frequencies as well as the corresponding resonance radii for frequency ratios of 1:1, 3:2, 4:3, and 5:4. Our results show that increasing $\alpha$ leads to deviations from the Schwarzschild case in both upper and lower QPO frequencies correlations and QPO orbital radii, with model-dependent trends. Further, we constrain the BH parameters using the observational data using MCMC analysis. Finally, we calculate the upper and lower QPO frequencies for a few BH candidates on the basis of the RP model using the constrained parameter values and find a good agreement with the observed results.
@@ I Introduction
One of the significant quantum effects that naturally emerges in the EFT framework is the trace anomaly, also known as the conformal anomaly [56, 57]. The classical field theories are conformally invariant and they exhibit a vanishing trace of the energy-momentum tensor. This symmetry is generally broken at the quantum level. In curved spacetime, this anomaly leads to a non-zero expectation value of the trace of the energy-momentum tensor. The trace anomaly captures the effects of quantum fields on the curvature of spacetime by adding specific geometric terms to the effective action. These terms depend on how spacetime is curved and include quantities like the square of the Weyl tensor, which measures the tidal distortion of spacetime, and the Gauss-Bonnet (GB) invariant, which combines different curvature components into a topological quantity. Together, these curvature dependent terms reflect how quantum corrections modify the classical description of gravity. These anomaly-induced corrections have profound implications for the semiclassical dynamics of spacetime, influencing phenomena like BH QNMs, [38, 39], quasi-periodic oscillations (QPOs) of BH [58, 59, 60, 61, 62], BH evaporation [40, 41, 42], quantum backreaction [43, 44], and early universe cosmology [45, 46, 47] etc. Within the EFT framework, the trace anomaly provides a concrete and calculable imprint of quantum effects in gravity, enabling theoretical predictions that remain compatible with the classical limit of GR. The effects of quantum corrections from EFT on gravity are typically very weak. However, they become significant in regions of extreme gravity, such as near the event horizon of a BH as mentioned earlier.
@@ II Black Hole Solutions in Trace Anomaly-Induced Gravity
$S=\int\mathrm{d}^{4}x\sqrt{-\bar{g}}\left[M^{2}e^{2\phi}\left(\bar{R}-12\bar{X}\right)-\bar{M}^{2}\bar{R}-\alpha\left(\phi\bar{\mathcal{G}}-4\,\bar{G}^{\mu\nu}\bar{\nabla}_{\mu}\phi\,\bar{\nabla}_{\nu}\phi+8\,\bar{X}\bar{\Box}\phi-8,\bar{X}^{2}\right)\right]$ (6)
@@
$\frac{\bar{h}^{\prime}}{\bar{h}}$ $-\frac{\bar{f}^{\prime}}{f}=\frac{2(\phi^{\prime 2}-\phi^{\prime\prime})\left[M^{2}\bar{r}^{2}e^{2\phi}+2\alpha\left(\bar{r}^{2}\bar{h}\phi^{\prime 2}+2\bar{r}\bar{h}\phi^{\prime}+\bar{h}-1\right)\right]}{\bar{r}M^{2}e^{2\phi}(\bar{r}\phi^{\prime}+1)-\bar{r}\bar{M}^{2}+2\alpha\phi^{\prime}\left(\bar{r}^{2}\bar{h}\phi^{\prime 2}+3\bar{r}\bar{h}\phi^{\prime}+3\bar{h}-1\right)}.$ (11)
@@
Here, $\bar{f}_{i}(\bar{r})$, $\bar{h}_{i}(\bar{r})$ and $\phi_{i}(\bar{r})$ denote functions of the radial coordinate $\bar{r}$, and $\bar{r}_{h}=2\mathcal{M}$ represents the event horizon radius of the BH, with $\mathcal{M}$ being its mass. For validity of Eq. (15) requires the following condition to be fulfilled:
@@
Thus, Eqs. (20) and (21) represent the black hole solutions in the presence of the GB trace anomaly, accurate up to second order in $\alpha$. In the following sections, we will make use of these solutions for further investigations.
@@ III Particle Motion Around Black Holes Modified by Trace Anomaly
In this section, we investigate the motion of test particles under different conditions in the spacetime of BHs influenced by the GB trace anomaly.
@@ B. Circular Orbits
Fig. 2 depicts the radial dependence of specific angular momentum and energy of the particle. The left panel shows the variation of the specific angular momentum with respect to $r\mathcal{M}^{-1}$. With the increase in the value of $\alpha$, the minimum of the angular momentum shifts to lower values compared to the Schwarzschild BH. The middle panel shows the variation of the energy $\mathcal{E}$, with respect to $r\mathcal{M}^{-1}$. In this case also with an increase in the value of the GB trace anomaly coupling constant $\alpha$, the minimum of the energy shifts to lower values compared to the Schwarzschild BH. The minimum in the energy profile corresponds to the most stable circular orbit. The right panel depicts the variation of the specific energy $\mathcal{E}$ with specific angular momentum $\mathcal{J}$. As $\alpha$ increases, the curves shift, leading to a decrease in the required angular momentum for a given energy. Compared to the Schwarzschild case, the presence of the trace anomaly allows circular orbits with lower angular momentum for the same energy, highlighting the influence of the GB trace anomaly on particle dynamics near BHs.
@@ A. Keplerian Frequencies
It is evident from Fig. 4 that the magnitude of the Keplerian frequency decreases monotonically with increasing $r\mathcal{M}^{-1}$ as expected for circular orbits around a BH. It should be noted that the presence of the GB term introduces modifications to the frequency profile. Specifically, larger values of the coupling constant $\alpha$ correspond to slightly higher frequencies at a given radius when compared to the Schwarzschild case, highlighting the influence of the GB correction on the orbital dynamics.
@@ B. QPO Orbits
$a\nu_{U}\left(\mathcal{M},r,l\right)=b\nu_{L}\left(\mathcal{M},r,l\right),$ (36)
@@ Appendix A Expressions of $\mathcal{J}$, $\mathcal{E}$, $\Omega_{r}$, $\Omega_{\theta}$ and $\Omega_{\phi}$
$\Omega_{r}=\sqrt{\frac{(2m-r)}{518400\,m^{3}M^{4}r^{8}\bar{M}^{2}}\left(\begin{aligned} &144000\alpha^{2}m^{5}\bar{M}^{2}+172800\alpha m^{4}M^{2}r^{3}\bar{M}^{2}+215872\alpha^{2}m^{4}r\bar{M}^{2}+259200m^{3}M^{4}r^{6}\bar{M}^{2}\\ &+158400\alpha m^{3}M^{2}r^{4}\bar{M}^{2}+267032\alpha^{2}m^{3}r^{2}\bar{M}^{2}+165600\alpha m^{2}M^{2}r^{5}\bar{M}^{2}+113956\alpha^{2}m^{2}r^{3}\bar{M}^{2}\\ &+53078\alpha^{2}mr^{4}\bar{M}^{2}+6979\alpha^{2}r^{5}\bar{M}^{2}+297600\alpha^{2}m^{5}M^{2}+89152\alpha^{2}m^{4}M^{2}r\\ &+6152\alpha^{2}m^{3}M^{2}r^{2}-90884\alpha^{2}m^{2}M^{2}r^{3}-52742\alpha^{2}mM^{2}r^{4}-38731\alpha^{2}M^{2}r^{5}\end{aligned}\right)}$ (45)
$\Omega_{\theta}=\Omega_{\phi}=\frac{1}{720}\sqrt{\frac{\begin{aligned} &2016000\,\alpha^{2}m^{6}\bar{M}^{2}+1382400\,\alpha\,m^{5}M^{2}r^{3}\bar{M}^{2}+1726464\,\alpha^{2}m^{5}r\,\bar{M}^{2}\\ &+518400\,m^{4}M^{4}r^{6}\bar{M}^{2}+432000\,\alpha\,m^{4}M^{2}r^{4}\bar{M}^{2}+1590960\,\alpha^{2}m^{4}r^{2}\bar{M}^{2}\\ &+345600\,\alpha\,m^{3}M^{2}r^{5}\bar{M}^{2}-156480\,\alpha^{2}m^{3}r^{3}\bar{M}^{2}\\ &-165600\,\alpha\,m^{2}M^{2}r^{6}\bar{M}^{2}-23400\,\alpha^{2}m^{2}r^{4}\bar{M}^{2}-78240\,\alpha^{2}m\,r^{5}\bar{M}^{2}\\ &-6979\,\alpha^{2}r^{6}\bar{M}^{2}+4166400\,\alpha^{2}m^{6}M^{2}-715776\,\alpha^{2}m^{5}M^{2}r\\ &-384240\,\alpha^{2}m^{4}M^{2}r^{2}-751680\,\alpha^{2}m^{3}M^{2}r^{3}-43800\,\alpha^{2}m^{2}M^{2}r^{4}\\ &-49440\,\alpha^{2}mM^{2}r^{5}+38731\,\alpha^{2}M^{2}r^{6}\end{aligned}}{m^{3}M^{4}r^{9}\bar{M}^{2}}}$ (46)

### v2

# Black Hole Quasi-Periodic Oscillations in the Presence of Gauss–Bonnet Trace Anomaly

@@ Abstract
We investigate the effects of the Gauss–Bonnet (GB) gravitational trace anomaly on the circular motion of test particles around black holes (BHs) and its implications for quasi-periodic oscillations (QPOs) in various theoretical models. Beginning with the equations of motion, we study the effects on effective potential, angular momentum, specific energy, and the innermost stable circular orbit (ISCO) induced by the anomaly parameter $\alpha$. The fundamental frequencies are calculated. Moreover, we examine several QPO models, including PR, RP, WD, TD, and ER2–ER4, and study the relationship between the upper and lower QPO frequencies as well as the corresponding resonance radii for frequency ratios of 1:1, 3:2, 4:3, and 5:4. Our results show that increasing $\alpha$ leads to deviations from the Schwarzschild case in both upper and lower QPO frequencies correlations and QPO orbital radii, with model-dependent trends. Further, we constrain the BH parameters using the observational data using MCMC analysis. Finally, we calculate the upper and lower QPO frequencies for a few BH candidates on the basis of the RP model using the constrained parameter values and find a good agreement with the observed results.
@@ I Introduction
One of the significant quantum effects that naturally emerges in the EFT framework is the trace anomaly, also known as the conformal anomaly [56, 57, 58]. The classical field theories are conformally invariant and they exhibit a vanishing trace of the energy-momentum tensor. This symmetry is generally broken at the quantum level. In curved spacetime, this anomaly leads to a non-zero expectation value of the trace of the energy-momentum tensor. The trace anomaly captures the effects of quantum fields on the curvature of spacetime by adding specific geometric terms to the effective action. These terms depend on how spacetime is curved and include quantities like the square of the Weyl tensor, which measures the tidal distortion of spacetime, and the Gauss-Bonnet (GB) invariant, which combines different curvature components into a topological quantity. Together, these curvature dependent terms reflect how quantum corrections modify the classical description of gravity. These anomaly-induced corrections have profound implications for the semiclassical dynamics of spacetime, influencing phenomena like BH QNMs, [38, 39], quasi-periodic oscillations (QPOs) of BH [59, 60, 61, 62, 63, 64, 65], BH evaporation [40, 41, 42], quantum backreaction [43, 44], and early universe cosmology [45, 46, 47] etc. Within the EFT framework, the trace anomaly provides a concrete and calculable imprint of quantum effects in gravity, enabling theoretical predictions that remain compatible with the classical limit of GR. The effects of quantum corrections from EFT on gravity are typically very weak. However, they become significant in regions of extreme gravity, such as near the event horizon of a BH as mentioned earlier.
@@ II Black Hole Solutions in Trace Anomaly-Induced Gravity
$S=\int\mathrm{d}^{4}x\sqrt{-\bar{g}}\left[M^{2}e^{2\phi}\left(\bar{R}-12\bar{X}\right)-\bar{M}^{2}\bar{R}-\alpha\left(\phi\bar{\mathcal{G}}-4\,\bar{G}^{\mu\nu}\bar{\nabla}_{\mu}\phi\,\bar{\nabla}_{\nu}\phi+8\,\bar{X}\bar{\Box}\phi-8\bar{X}^{2}\right)\right].$ (6)
@@
$\frac{\bar{h}^{\prime}}{\bar{h}}$ $-\frac{\bar{f}^{\prime}}{\bar{f}}=\frac{2(\phi^{\prime 2}-\phi^{\prime\prime})\left[M^{2}\bar{r}^{2}e^{2\phi}+2\alpha\left(\bar{r}^{2}\bar{h}\phi^{\prime 2}+2\bar{r}\bar{h}\phi^{\prime}+\bar{h}-1\right)\right]}{\bar{r}M^{2}e^{2\phi}(\bar{r}\phi^{\prime}+1)-\bar{r}\bar{M}^{2}+2\alpha\phi^{\prime}\left(\bar{r}^{2}\bar{h}\phi^{\prime 2}+3\bar{r}\bar{h}\phi^{\prime}+3\bar{h}-1\right)}.$ (11)
@@
Here, $\bar{f}_{i}(\bar{r})$, $\bar{h}_{i}(\bar{r})$ and $\phi_{i}(\bar{r})$ denote functions of the radial coordinate $\bar{r}$, and $\bar{r}_{h}$ represents the event horizon radius of the BH in the barred frame. However, for consistency with the subsequent analysis performed in the unbarred coordinate system defined by $g_{\mu\nu}$, i.e., in the unbarred frame, we identify the physical event-horizon radius as $r_{h}=2\mathcal{M}$, with $\mathcal{M}$ being the mass of the BH. For validity of Eq. (15) requires the following condition to be fulfilled:
@@
Thus, Eqs. (20) and (21) represent the black hole solutions in the presence of the GB trace anomaly, accurate up to second order in $\alpha$. In the following sections, we will make use of these solutions for further investigations. It is to be noted here that in the underlying quantum theory, the coupling parameter $\alpha$ is proportional to the GB anomaly coefficient of the conformal matter field, which is non-negative for any unitary four-dimensional conformal field theory and is therefore non-negative in physically consistent anomaly-induced effective actions [78, 77, 79]. For this reason, and following the conventions of anomaly-induced effective actions, we restrict our analysis to $\alpha\geq 0$ in the present work.
@@ III Particle Motion Around Black Holes Modified by Trace Anomaly
In this section, we investigate the motion of test particles under different conditions in the spacetime of BHs influenced by the GB trace anomaly. In all numerical computations, we work in geometrized units, $G=c=1$. Following the standard practice in BH accretion and QPO studies, we express all radii and frequencies in units of the BH mass $\mathcal{M}$, so that these quantities become dimensionless. In practice, we set $\mathcal{M}=1$ for all numerical calculations. Moreover, in the unbarred frame, the perturbative condition (16) takes the form $\alpha/M^{2}r_{h}^{2}$, where $M$ is the new mass scale appearing in the anomaly-induced action, related to the reduced Planck mass by $M=M_{\rm pl}/\sqrt{2}$. In our analysis, we consider values of the anomaly parameter $\alpha$ of the order of $10^{-5}$, and we identify the horizon radius as $r_{h}=2\mathcal{M}$. Thus in the chosen unit of $\mathcal{M}=1$ along with $M_{\rm pl}=1$, the value of $\alpha/M^{2}r_{h}^{2}$ become the order of $\sim 10^{-5}$ in the explored parameter range. Since this is several orders of magnitude smaller than $1$, the perturbative condition (16) will be safely satisfied throughout the entire parameter space used in this work.
@@ B. Circular Orbits
In Fig. 2, we plot the specific energy $\mathcal{E}$ and specific angular momentum $\mathcal{J}$ (not their squared values). Both quantities $\mathcal{E}$ and $\mathcal{J}$ are in the unit normalised by the BH mass $\mathcal{M}$. This figure depicts the radial dependence of specific angular momentum and energy of the particle. The left panel shows the variation of the specific angular momentum with respect to $r\mathcal{M}^{-1}$. With the increase in the value of $\alpha$, the minimum of the angular momentum shifts to lower values compared to the Schwarzschild BH. The middle panel shows the variation of the energy $\mathcal{E}$, with respect to $r\mathcal{M}^{-1}$. In this case also with an increase in the value of the GB trace anomaly coupling constant $\alpha$, the minimum of the energy shifts to lower values compared to the Schwarzschild BH. The minimum in the energy profile corresponds to the most stable circular orbit. The right panel depicts the variation of the specific energy $\mathcal{E}$ with specific angular momentum $\mathcal{J}$. As $\alpha$ increases, the curves shift, leading to a decrease in the required angular momentum for a given energy. Compared to the Schwarzschild case, the presence of the trace anomaly allows circular orbits with lower angular momentum for the same energy, highlighting the influence of the GB trace anomaly on particle dynamics near BHs.
@@ A. Keplerian Frequencies
It is evident from Fig. 4 that the magnitude of the Keplerian frequency $\Omega_{\phi}\mathcal{M}^{-1}$ decreases monotonically with increasing $r\mathcal{M}^{-1}$, as expected for circular orbits around a BH. When the GB trace anomaly correction is included, the resulting frequency profiles remain very close to the Schwarzschild curve for $r\mathcal{M}^{-1}>r_{h}\mathcal{M}^{-1}$ ($r_{h}\mathcal{M}^{-1}=2$ in our case). This behaviour reflects the fact that the anomaly induced modification to the metric is a small perturbation that falls off rapidly with radius.
@@ B. QPO Orbits
$a\nu_{U}\left(\mathcal{M},r,\alpha\right)=b\nu_{L}\left(\mathcal{M},r,\alpha\right),$ (36)
@@ Appendix A Expressions of $\mathcal{J}$, $\mathcal{E}$, $\Omega_{r}$, $\Omega_{\theta}$ and $\Omega_{\phi}$
$\Omega_{r}=\sqrt{\frac{(2\mathcal{M}-r)}{518400\,\mathcal{M}^{3}M^{4}r^{8}\bar{M}^{2}}\left(\begin{aligned} &144000\alpha^{2}\mathcal{M}^{5}\bar{M}^{2}+172800\alpha\mathcal{M}^{4}M^{2}r^{3}\bar{M}^{2}+215872\alpha^{2}\mathcal{M}^{4}r\bar{M}^{2}+259200\mathcal{M}^{3}M^{4}r^{6}\bar{M}^{2}\\ &+158400\alpha\mathcal{M}^{3}M^{2}r^{4}\bar{M}^{2}+267032\alpha^{2}\mathcal{M}^{3}r^{2}\bar{M}^{2}+165600\alpha\mathcal{M}^{2}M^{2}r^{5}\bar{M}^{2}+113956\alpha^{2}\mathcal{M}^{2}r^{3}\bar{M}^{2}\\ &+53078\alpha^{2}\mathcal{M}r^{4}\bar{M}^{2}+6979\alpha^{2}r^{5}\bar{M}^{2}+297600\alpha^{2}\mathcal{M}^{5}M^{2}+89152\alpha^{2}\mathcal{M}^{4}M^{2}r\\ &+6152\alpha^{2}\mathcal{M}^{3}M^{2}r^{2}-90884\alpha^{2}\mathcal{M}^{2}M^{2}r^{3}-52742\alpha^{2}\mathcal{M}M^{2}r^{4}-38731\alpha^{2}M^{2}r^{5}\end{aligned}\right)}$ (45)
$\Omega_{\theta}=\Omega_{\phi}=\frac{1}{720}\sqrt{\frac{\begin{aligned} &2016000\,\alpha^{2}\mathcal{M}^{6}\bar{M}^{2}+1382400\,\alpha\,\mathcal{M}^{5}M^{2}r^{3}\bar{M}^{2}+1726464\,\alpha^{2}\mathcal{M}^{5}r\,\bar{M}^{2}\\ &+518400\,\mathcal{M}^{4}M^{4}r^{6}\bar{M}^{2}+432000\,\alpha\,\mathcal{M}^{4}M^{2}r^{4}\bar{M}^{2}+1590960\,\alpha^{2}\mathcal{M}^{4}r^{2}\bar{M}^{2}\\ &+345600\,\alpha\,\mathcal{M}^{3}M^{2}r^{5}\bar{M}^{2}-156480\,\alpha^{2}\mathcal{M}^{3}r^{3}\bar{M}^{2}\\ &-165600\,\alpha\,\mathcal{M}^{2}M^{2}r^{6}\bar{M}^{2}-23400\,\alpha^{2}\mathcal{M}^{2}r^{4}\bar{M}^{2}-78240\,\alpha^{2}\mathcal{M}\,r^{5}\bar{M}^{2}\\ &-6979\,\alpha^{2}r^{6}\bar{M}^{2}+4166400\,\alpha^{2}\mathcal{M}^{6}M^{2}-715776\,\alpha^{2}\mathcal{M}^{5}M^{2}r\\ &-384240\,\alpha^{2}\mathcal{M}^{4}M^{2}r^{2}-751680\,\alpha^{2}\mathcal{M}^{3}M^{2}r^{3}-43800\,\alpha^{2}\mathcal{M}^{2}M^{2}r^{4}\\ &-49440\,\alpha^{2}\mathcal{M}M^{2}r^{5}+38731\,\alpha^{2}M^{2}r^{6}\end{aligned}}{\mathcal{M}^{3}M^{4}r^{9}\bar{M}^{2}}}$ (46)